### v3

# Adversarial Coevolutionary Illumination with Generational Adversarial MAP-Elites

Abstract
Quality-Diversity (QD) algorithms seek to discover diverse, high-performing solutions across a behavior space, contrasting with conventional optimization methods that target a single optimum. Adversarial problems present unique challenges for QD approaches, as the competing nature of opposing sides creates interdependencies that complicate the evolution process. Existing QD methods applied to such scenarios typically fix one side, constraining behavioral diversity. We present Generational Adversarial MAP-Elites (GAME), a coevolutionary QD algorithm that evolves both sides by alternating which side is evolved at each generation. By integrating a vision embedding model, our approach eliminates the need for domain-specific behavior descriptors and instead operates on video. We validate GAME across three distinct adversarial domains: a multi-agent battle game, a soft-robot wrestling environment, and a deck building game. Our experiments reveal several evolutionary phenomena, including arms-race-like dynamics, enhanced novelty through generational extinction, and the preservation of neutral mutations as crucial stepping stones toward the highest performance. While GAME successfully illuminates all adversarial problems, its capacity for truly open-ended discovery remains constrained by the finite nature of the underlying search spaces. These findings establish GAME’s broad applicability while highlighting opportunities for future research into open-ended adversarial coevolution.
Code and videos available at: https://github.com/Timothee-ANNE/GAME
Introduction
Figure 1: GAME’s illumination of a multi-agent adversarial game. The point cloud is a 2D PCA projection (2222% and 1212% explained variance) of the intergenerational tournament between elites found for one run of GAME across 2020 generations with 100 000100\,000 evaluations per generation. We display timed snapshots of eight duels exhibiting different behaviors. We also indicate the fitness of both sides, represented as the percentage of the opposing side’s depleted health. Videos of these duels and supplementary data are available in the repository https://github.com/Timothee-ANNE/GAME.
Quality Diversity (QD) (Pugh et al.,, 2016) is a branch of evolutionary algorithms that seeks to illuminate a problem, i.e., to discover a diverse set of high-quality solutions that maximize a fitness function while covering a behavior space. It has demonstrated success across various domains, including robotics (Cully et al.,, 2015), procedural content generation (Gravina et al.,, 2019), chemical synthesis (Jiang et al.,, 2022), and aeronautics (Brevault and Balesdent,, 2024).
One may wish to apply such methods to adversarial problems, which arise across many domains, for example, in military conflicts (Schelling,, 1980), economic regulation (Baldwin et al.,, 2011), machine learning (Chakraborty et al.,, 2018), and cybersecurity (Li et al.,, 2021).
Several such efforts have already been undertaken. For instance, in the turn-based strategy card game Hearthstone (Blizzard Entertainment,, 2014), Fontaine et al., (2019) illuminate the deck space, while Fontaine et al., (2020) evolve a set of neural network controllers with a fixed deck to find different strategies. More recently, Samvelyan et al., 2024a investigate strategic weaknesses in a Deep Reinforcement Learning (DRL) agent during a simulated football game, while Samvelyan et al., 2024b search for different safety-violating prompts for a Large Language Model (LLM). A limitation of these approaches is that they only optimize one side of the adversarial problem. This is problematic, as adversarial systems often lead to an arms race (Dawkins and Krebs,, 1979), which requires improvements on both sides to emerge.
Optimizing both sides simultaneously to broaden the illumination aligns with the artificial life challenge of fostering open-ended evolution through adversarial coevolution (Bedau et al.,, 2000; Dorin and Stepney,, 2024). Some works explore this avenue: Costa et al., (2020) coevolve Generative Adversarial Networks (GANs) using QD, leading to improved models; Wang et al., (2019, 2020) coevolve bipedal walkers and their environments to foster open-ended evolution, creating an automatic curriculum; and Dharna et al., (2024) introduce a QD algorithm that generates Python scripts to control both sides of a car chase game. However, these works rely on problem-specific methods.
This paper significantly extends our conference paper (Anne et al., 2025a, ), providing a comprehensive introduction to Generational Adversarial MAP-Elites (GAME), a new QD algorithm for adversarial problems, applied to three different adversarial domains: (1) a multi-agent battle game called Parabellum (Anne et al., 2025b, ); (2) a 2D soft-robot custom EvoGym called Wrestling (Bhatia et al.,, 2021); and (3) a deck building game called Hearthbreaker (Daniel et al.,, 2014), which is a Python simulator of Hearthstone (Blizzard Entertainment,, 2014).
Combined with a Vision Embedding Model (VEM) such as CLIP (Radford et al.,, 2021), GAME can operate directly on videos of behaviors, eliminating the need for handcrafted behavior descriptors. GAME handles the high dimensionality of the embedding space by using growing unstructured archives (Vassiliades et al.,, 2017).
We demonstrate through ablation studies in Parabellum that all components of GAME are necessary to illuminate the space of possible strategies effectively. In Wrestling, we show that GAME shows promise for the evolution of artificial creatures, and in Hearthbreaker, we show that GAME’s adversarial coevolution leads to solutions of higher quality but smaller coverage than applying MAP-Elites against a fixed opponent deck.
The paper’s main contributions are:
  • a new QD algorithm, GAME, for adversarial problems illumination;
  • the use of GAME with a VEM as a domain-agnostic behavior space in two distinct domains, a multi-agent battle game and a 2D soft-robot wrestling environment, showing GAME’s generality;
  • the comparison of GAME (without the VEM) against a one-sided illumination in a deck building domain.
Related Work
Artificial Life, Open-endedness, and Coevolution
GAME aims to create an open-ended adversarial coevolution of solutions that broadens illumination. Creating an open-ended artificial environment that continually presents new, interesting challenges and, in turn, fosters the evolution of novel and meaningful solutions to those challenges is one of the main quests in the artificial life community (Bedau et al.,, 2000; Dorin and Stepney,, 2024).
A way to move toward open-endedness is through coevolution in an arms race dynamic (Dawkins and Krebs,, 1979). A key difficulty lies in finding a balance that avoids the collapse of one side or the emergence of a stable equilibrium where both sides cease to innovate (Ficici and Pollack,, 1998). Moran and Pollack, (2019) demonstrate that a blend of cooperation and competition can lead to greater complexity growth. Similarly, Harrington and Pollack, (2019) show that greater policy evolvability supports a longer-lasting arms race and increased complexity.
A form of not strictly adversarial problem is the coevolution of an environment and the agent interacting with it. The objective is for the environment to evolve in a way that presents new challenges that are neither too easy nor too difficult, effectively creating a curriculum for the agent by offering a sequence of meaningful stepping stones. Brant and Stanley, (2017) coevolve mazes and maze-solving agents using minimal criterion coevolution, i.e., any agent that solves a maze and any maze that is solved at least once can reproduce. Their approach demonstrates that no handcrafted fitness function or behavior descriptor is required.
Building on this idea, POET (Wang et al.,, 2019) generates an open-ended sequence of environments for training bipedal walkers. It maintains a population of active environment–controller pairs. New environments are generated from recently active ones that have shown sufficient progress where the task is neither easy nor difficult for the associated controller. Each controller is independently optimized using an evolutionary strategy. POET also facilitates transfer between pairs, attempting to apply controllers trained in one environment to others. A key insight is that some resulting controllers could not be obtained through direct optimization on the final environment, suggesting that POET enables the discovery of critical stepping stones.
Enhanced-POET (Wang et al.,, 2020) introduces a new measure of novelty for environments, called Performance of All Transferred Agents – Environment Comparison (PATA-EC). This metric requires only fitness scores, eliminating the need for handcrafted behavior descriptors. It is based on a round-robin tournament of agents, comparing their performance rankings across environments. The underlying idea is that a novel environment should yield a different ranking.
Those works are mainly designed to propose an automatic curriculum learning to discover robust agents that can solve challenging environments without explicitly trying to illuminate the space of agents or environments. In contrast, GAME explicitly aims to illuminate both sides of the adversarial problem.
Self-Play
Self-play is a closely related field, which creates a learning curriculum by making the agent play against itself or an earlier version of itself. AlphaStar (Vinyals et al.,, 2019) achieves grandmaster level in the real-time strategy (RTS) player-versus-player game StarCraft 2. It uses a league of agents that serve as training opponents for the main agents. Using a league prevents strategy collapse, improves robustness, and helps avoid local minima and forgetting.
Baker et al., (2019) propose an RL framework for adversarial self-play in a 3D multi-agent hide-and-seek game. They revealed how it allowed the emergence of six strategy phases as both sides found new ways to counter the other’s high-performing behaviors. The goal of self-play is, however, the learning of one robust and high-performing policy, not the illumination of all possible strategies.
Quality-Diversity for adversarial problems
Two popular QD algorithms are Novelty Search with Local Competition (NSLC) (Lehman and Stanley,, 2011) and Multidimensional Archive of Phenotypic Elites (MAP-Elites) (Mouret and Clune,, 2015). MAP-Elites works by constructing an archive of high-performing solutions, known as elites, which are organized into different cells that discretize the behavior space. MAP-Elites generates a new candidate solution at each iteration using evolutionary variation operators, evaluates it, and if it either exhibits a novel behavior (i.e., fills a previously empty behavior cell) or has a greater fitness than the current elite in its corresponding cell, it becomes the elite of that cell.
Multi-Task MAP-Elites (MT-ME) (Mouret and Maguire,, 2020) is a variant of MAP-Elites that addresses multi-task problems, i.e., the simultaneous optimization of multiple fitness functions. The idea is that solving these tasks simultaneously can improve sampling efficiency, as related tasks may share similar solutions. Building on this, Multi-task Multi-behavior MAP-Elites (MTMB-ME) (Anne and Mouret,, 2023) extends MT-ME by evolving diverse high-performing solutions for each task, rather than just a single solution. In this paper, GAME uses MTMB-ME at each generation to evolve solutions that compete against a set of solutions from previous generations (i.e., the tasks).
QD algorithms have been used to illuminate adversarial problems. For example, Fontaine et al., (2019) propose MAP-Elites with sliding boundaries to build different decks in Hearthstone (Blizzard Entertainment,, 2014), using as opponents first a fixed set of decks and then the best decks found in the first phase to explore the possible counters. Using the six best found decks as opponents, Fontaine et al., (2020) use Covariance Matrix Adaptation MAP-Elites to evolve a diverse set of neural network controllers and find different strategies when playing with one deck. Steckel and Schrum, (2021) use MAP-Elites to illuminate the space for generated levels by different GANs in the Lone Runner game. Wan et al., (2024) apply QD principles to adversarial imitation learning to train agents on several behaviors simultaneously and show that it can potentially improve any inverse RL method. Samvelyan et al., 2024a apply MAP-Elites to explore adversarial scenarios for a pre-trained DRL agent in a multi-agent game environment (Google Research Football), uncovering strategic weaknesses in the agent’s behavior. Samvelyan et al., 2024b use MT-ME to find a set of diverse adversarial safety attacks on an LLM, using an LLM for the variation operator and for evaluating the fitness of the attack. These methods only consider optimizing one side of the adversarial problem, significantly limiting the illumination. The third case study, Hearthbreaker, is an initial attempt to compare GAME’s ability to alternate the illumination against a one-sided illumination similar to Fontaine et al., (2019). A noteworthy difference is that we use the Hearthbreaker (Daniel et al.,, 2014) simulator in Python, while Fontaine et al., (2019) use the SabberStone (Milva et al.,, 2016) simulator in C#, which proposes better AI agents for playing the deck.
Few QD methods currently try to coevolve both sides. Costa et al., (2020) use NSLC to coevolve GANs (i.e., the generators and descriptors), showing that it improves diversity and discovers better models. Closer to our work, Dharna et al., (2024) propose a self-play QD algorithm to coevolve adversarial controllers as Python code for an asymmetric adversarial game between pursuer and evader agents. They also use the embedding space of a foundation model for diversity, but use an LLM embedding space of the Python controller script, which relates to genotypic diversity rather than behavioral diversity, as in this paper. Their work also uses an unstructured archive to handle the high-dimensional embedding space. However, they employ an NSLC-type unstructured archive to define novelty, which requires more problem-specific tuning of the novelty threshold parameter compared to the growing archives used in GAME.
Using a VEM for behavior space
Figure 2: GAME is an adversarial coevolution QD algorithm that iterates the illumination of one side of an adversarial problem using MTMB-ME (Anne and Mouret,, 2023), switching sides at each generation to promote arms-race dynamics that expand the illumination. One key feature of GAME is the ability to use a VEM as a behavior space.
One limitation of MAP-Elites is the need to define a behavior space to determine what is different. Automated Search for Artificial Life (ASAL) (Kumar et al.,, 2024) has shown that a VEM, such as CLIP (Radford et al.,, 2021), can be used to explore the space of artificial life substrates. Inspired by them, we propose, to our knowledge, the first use of a VEM as a behavior space for a MAP-Elites algorithm.
One difficulty with such a space is its high dimensionality (e.g., 2 5602\,560 dimensions in this paper). The original MAP-Elites uses a grid to divide the behavior space, which is not scalable to high dimensions (as the number of cells grows exponentially with the dimensions). One solution is using Centroidal Voronoi Tessellations (CVT) (Vassiliades et al.,, 2016) to divide the behavior space uniformly. However, in very high-dimensional spaces, one cannot easily predict what proportion of the space will be covered by the studied system, making it difficult to tune the selection pressure. If too many cells are available to fill, the pressure will be too low, reducing quality, and if there are not enough cells, the pressure will be too high, reducing diversity. One solution is to learn a low-dimensional representation using an autoencoder (Cully,, 2019), which requires extra computation.
Our intuition is that we can grow the archive to cover the relevant part of the behavior space instead of defining it beforehand. This is similar to how NSLC defines novelty with a growing archive of solutions. This is called an unstructured archive (Vassiliades et al.,, 2017). Unlike NSLC, which uses a distance-threshold parameter to determine whether a new solution should be added to the archive, unstructured archives have a predefined maximal number of cells and only update the position and boundaries of those cells. This can be done by updating the centroids of the CVT when the replacement of an old centroid by a new one would increase the minimal distance between all centroids, thus growing the archive.
Generational Adversarial MAP-Elites
GAME is a QD algorithm that aims to illuminate both sides (Blue and Red) of an adversarial problem using coevolution to nurture open-ended discovery of solutions (Fig. 2). It is divided into two levels: (1) intergenerational illumination of solutions, switching from one side to the other at each generation, Alg. 1 (a, f, and g); and (2) intragenerational execution of MTMB-ME at each generation, using one side as tasks and the other as solutions, Alg. 2 (b, c, d, and e).
It proceeds as follows (the paper’s values in parentheses):
  • (a)
    randomly sample $N_{task}$ ($100$) Blue solutions as tasks;
    For $N_{gen}$ (2020) generations:
    • (b)
      initialize a multi-task multi-behavior growing archive with $N_{cells}$ ($25$) for each task;
      For $N_{budget}$ (100 000100\,000) evaluations:
      • (c)
        pick a tasktask at random and use a variation operator on randomly picked elites from the whole archive to generate a candidate solution $s$;
      • (d)
        evaluate $s$ against the tasktask, collect the video and fitness $f$, subsample the video (five uniformly spaced frames) to get framesframes, obtain the embedding (512512D) from each using a VEM (CLIP), and concatenate to get the behavior descriptor $b$ (2 5602\,560D);
      • (e)
        update tasktask’s archive (Alg. 3): (1) if the new behavior $b$ is farther from all the current cells’ centroids than the closest pair of centroids, it is added to the archive and one of the two centroids from the pair is removed; (2) else if the new fitness $f$ is greater than the corresponding cell’s fitness, then the new solution becomes the elite of its cell;
    • (f)
      aggregate all the elites, cluster them into $N_{tasks}$ clusters using K-Means, and select the elite of each cluster to form the next generation of tasks;
    • (g)
      bootstrap the next generation with a tournament between the new and previous tasks.
Algorithm 1 GAME
Inputs: Red search space $S_{Red}$, Blue search space $S_{Blue}$
Parameters: $N_{gen}$, $N_{task}$
1: $Tasks\leftarrow$ Sample $N_{task}$ random Blue solutions $\triangleright$ (a)
2: $Generations=\{0:Tasks\}$
3: $\mathcal{B}=\emptyset$ $\triangleright$ For storing bootstrapping evaluations
4:for gen_id=1gen\_id=1 in $N_{gen}$ do
5:   S $\leftarrow$ $S_{Red}$ if gen_idgen\_id is odd else $S_{Blue}$
6:   $\mathcal{A}\leftarrow\operatorname{MTMB-ME}(Tasks,S,\mathcal{B})$ $\triangleright$ (b-e) - Alg. 2
7:   BehaviorsBehaviors $\leftarrow$ Aggregate the archive’s elites $\triangleright$ (f)
8:   Clusters $\leftarrow\operatorname{K-means}(Behaviors,k=N_{task})$
9:   $Tasks\leftarrow\{\operatorname{Elite}(cluster)\}_{cluster\in Clusters}$
10:   Generations[gen_id]=TasksGenerations[gen\_id]=Tasks
11:   $\mathcal{B}\leftarrow$ TasksTasks versus Generations[gen_id1]Generations[gen\_id-1] $\triangleright$ (g)
12: return GenerationsGenerations
Algorithm 2 MTMB-ME with growing archive and a VEM
Inputs: $Tasks$, search space $S$, bootstrap set $\mathcal{B}$
Parameters: Number of evaluations $N_{budget}$, Number of initial random search $N_{init}$, evaluation function $\operatorname{Evaluate}$, variation operator $\operatorname{Variation}$, VEM
1: $\mathcal{A}\leftarrow$ Initialize $N_{task}$ growing archives $\triangleright$ (b)
2:for $(task,s,f,b)$ in $\mathcal{B}$ do $\triangleright$ (g) - Bootstrapping
3:   $\mathcal{A}\leftarrow\operatorname{Update}(\mathcal{A}[task],s,f,b)$ $\triangleright$ Alg. 3
4:for $i=1$ to $N_{budget}$ do $\triangleright$ Main loop
5:   $task\leftarrow$ Select a task at random from TasksTasks $\triangleright$ (c)
6:  if $\mathcal{A}$ has fewer than $N_{init}$ elites then
7:    $s\leftarrow$ Sample a random solution from $S$
8:  else
9:    $s\leftarrow\operatorname{Variation}(\mathcal{A})$
10:   $f$, $video\leftarrow\operatorname{Evaluate}(s,task)$ $\triangleright$ (d)
11:   $frames\leftarrow$ Subsample the videovideo
12:   $b$ $\leftarrow$ Apply the VEM on each frame
13:   $\mathcal{A}\leftarrow\operatorname{Update}(\mathcal{A}[task],s,f,b)$ $\triangleright$ (e) - Alg. 3
14: return Archives
Using a VEM for behavior space
When using a VEM, GAME follows ASAL by using cosine similarity to compute the distance between two behaviors (which are 2 5602\,560D):
$\operatorname{dist}(b,b^{\prime})=1-\dfrac{b\cdot b^{\prime}}{\|b\|_{2}\|b^{\prime}\|_{2}}.$
The reason is that Euclidean distance loses meaning as the number of dimensions increases. This does not impact the growing archive, as it uses relative comparisons of distances between behaviors rather than their actual positions.
Growing an unstructured archive
One point to consider when growing an archive is that when a new behavior takes over a cell, it can “steal” the elites of neighboring cells as it changes the whole CVT, which creates “holes.” A simple solution to fill those holes is to reinstate the centroid as the elite of its cell, as it can never be “stolen.” It only requires storing an additional solution per cell. More elaborate heuristics can be used, e.g., storing all the past elites and checking with the current CVT. However, we found that this does not significantly improve performance for most problems while increasing computation and memory usage.
Algorithm 3 Growing unstructured archive update
Inputs: archive $A$, solution $s$, fitness $f$, behavior $b$
Parameters: max archive size $N_{cells}$, distance function $\operatorname{dist}$
1: $(C,E,E_{backup})=A$ $\triangleright$ Centroids, Elites, and Backup Elites
2:if $\operatorname{size}(C)<N_{cells}$ then $\triangleright$ Add a new cell
3:   $i=\operatorname{size}(E)$
4:   $C_{i}=b$
5:   $E[i]\leftarrow(s,f,b)$
6:   $E_{backup}[i]\leftarrow(s,f,b)$
7:else $\triangleright$ Check behavior and fitness
8:   $distances=\{\operatorname{dist}(C_{i},C_{j})\}_{0\leq i<j<N_{cells}}$
9:   $d_{min}=\min(distances)$
10:   $d=\min\{\operatorname{dist}(b,C_{i})\}_{0\leq i<N_{cells}}$
11:   $c_{id}=\operatorname{find\_cell}(C,b)$ $\triangleright$ Closest centroid’s index
12:  if $d>d_{min}$ then $\triangleright$ New enough behavior = growth
13:    $j,k\leftarrow\operatorname{argmin}(distances)$
14:    $d_{j}\leftarrow\min\{\operatorname{dist}(C_{j},C_{i})\}_{0\leq i\neq j<N_{cells}}$
15:    $d_{k}\leftarrow\min\{\operatorname{dist}(C_{k},C_{i})\}_{0\leq i\neq k<N_{cells}}$
16:    $k\leftarrow j$ if $d_{j}<d_{k}$ else $k$
17:    $C_{k}\leftarrow b$
18:    $E[k]\leftarrow(s,f,b)$
19:    $E_{backup}[k]\leftarrow(s,f,b)$
20:   for $i=0$ to $N_{cells}$ do $\triangleright$ Check and repair holes
21:     if $\operatorname{find\_cell}(C,E[i].b)\neq i$ then
22:       $E[i]\leftarrow E_{backup}[i]$
23:  else if f >> E[$c_{id}$].f then $\triangleright$ Better fitness
24:    $E[c_{id}]\leftarrow(s,f,b)$
25: return $(C,E,E_{backup})$
Case study: Multi-agent adversarial game, Parabellum
Figure 3: GAME’s variants and ablations comparisons. The solid line represents the median, and the shaded area shows the min and max of 33 replications over 2020 generations. (a) Quality-only and Diversity-only show the largest increase in solution size, but (b) Quality-only and GAME-SO show the largest increase in complexity. (c–d) Quality-only and GAME-MO (no VEM) lead to the worst QD performances. (e) Removing bootstrapping leads to a constant discovery of novel solutions.
To evaluate GAME, we first applied it to an adversarial battle game research environment, Parabellum (Anne et al., 2025b, ). Battle games are inherently adversarial, making them a direct target for illuminating adversarial problems. In addition, battle games employ many agents in a top-down view of the map, and characterizing multi-agent behavior is not trivial. Parabellum also shares similarities with ALife simulations, where many agents move and interact on a 2D map, which inspired us to apply the same VEM as in ASAL (Kumar et al.,, 2024).
Parabellum Multi-Agent Game
Parabellum is a multi-agent battle game where two sides, Blue and Red, try to eliminate the other, i.e., the fitness of each side is the sum of the depleted health of the opposing side’s units. Each unit can only see other units within its sight range ($15$ m, with the map being $100$ m wide). At each time step, each unit can either move, stand, attack an enemy in reach, or heal an ally in reach, given its local observation (position, type, side, and health of all the units in sight). It is implemented in JAX (Bradbury et al.,, 2018), a Python library for just-in-time compilation and vectorization. It permits fast parallelization of the units’ behaviors using a dedicated GPU. Parabellum contains stochasticity, but we use the same seed for all encounters to make comparisons fair.
We defined five types of units (each with a different color for visualization): spearman: slow and high-health close combat unit; archer: low-health long-range combat unit; cavalry: fast and medium-health close combat unit; healer: low-health healing unit with close range; grenadier: low-health mid-range unit that inflicts area damage to all units.
Each side comprises $32$ units of each type uniformly spread in each side’s starting area, which are placed so that no unit initially sees an enemy unit (see Fig. 1 frames 1).
Behavior Tree as Controller
We pair Parabellum with behavior trees (BTs) (Colledanchise and Ögren,, 2018) to determine the behavior of each agent. BTs are commonly used in robotics and video games, intuitive to design, and inherently interpretable.
More specifically, the units from each side share the same BT, meaning that the search space of GAME is one BT for the Red side and one BT for the Blue side. At each evaluation, each unit visits its BT starting from the root node and undergoes a left-most depth-first visit of the tree until it finds a valid action. The BT comprises intermediate nodes: Sequence/Failwith nodes, which stop at the first invalid/valid node, and leaf nodes: Action nodes that return an action and its validity, and Condition nodes that return a boolean. To obtain fast evaluations, we implemented a BT evaluation function in JAX that allows vectorization of the BT evaluation at the price of setting a maximal number of leaves (100).
Their tree nature allows the use of genetic variation operators for trees. For example, Iovino et al., (2021) evolve BTs for robot control, and Montague et al., (2023) use MAP-Elites with BTs to learn controllers for heterogeneous robot swarms. Taking inspiration from them, GAME randomly selects a variation operator: deleting a subtree (3535%), adding a random node at a random location (2121%), mutating a node by changing its parameters (77%), replacing a node with a random node (77%), or performing a crossover, i.e., copying a random subtree of another BT at a random location (3030%). Those probabilities could be tuned but do not seem to change GAME’s performance significantly.
Another advantage of using BTs is that it is straightforward to increase their complexity by allowing them to grow, promoting open-ended evolution (Harrington and Pollack,, 2019). Such an increase in complexity is not trivial to obtain with neural networks and must rely on specific methods such as NEAT (Stanley and Miikkulainen,, 2002).
The BT evolution requires the design of atomic functions (the Actions and Conditions) that take the unit’s local observation and return the corresponding output. To allow synchronization between units, each side can set a target position on the map and move toward it even if it is out of sight. Different qualifiers parametrize the atomics to improve interpretation and usage, e.g., a target qualifier corresponding to either closest, farthest, weakest, strongest, or random.
The available Actions are: Stand (do nothing); Attack (attack the target enemy in reach of a given type); Heal (heal the target ally in reach of a given type); Move (move toward/away from the target ally/enemy of a given type); Go To (go to the target position); Set Target (mark the position of the target ally/enemy as target position).
The available Conditions are: In Sight (Is there an ally/enemy of a given type in sight?) In Reach (Is there an ally/enemy of a given type in reach?) Is Dying (Is my health or an ally’s/enemy’s health below a given threshold?) Is Type (Am I of a given type?) Is Set Target (Is the target position set on the map?)
GAME variants and ablations
To evaluate GAME, we evaluate six variants:
  • GAME-SO: full variant that uses a single-objective fitness to maximize the opposing side’s depleted health;
  • GAME-MO: full variant with a multi-objective fitness that first maximizes the opposing side’s depleted health and secondarily minimizes the size of the BT;
  • GAME-SO (no bootstrap): GAME-SO with an ablation of the bootstrapping between generations;
  • GAME-SO (no VEM): GAME-SO with a 2D handcrafted behavior space equal to the evaluated side’s average remaining health and the time before completion (i.e., one side wins or $100$ steps timeout) instead of a VEM;
  • Diversity-only: do not use fitness, only rule (e.1);
  • Quality-only: MT-ME instead of MTMB-ME, i.e., each task has only one solution.
We run three replications of $20$ generations for each variant with 100 000100\,000 evaluations of $100$ steps per generation, $100$ tasks per generation, and $25$ cells per task’s archive.
Measures
After each run, we conduct a tournament between each generation’s tasks, $10$ Blue generations against $10$ Red generations (i.e., 1 0001\,000 Red elites against 1 0001\,000 Blue elites), resulting in 1 000 0001\,000\,000 evaluations. This allows us to study whether there are intergenerational improvements.
Solution Size and Behavioral Complexity
To measure variations in solution size, we compute the number of nodes of the BT elites (Fig. 3.a), and for behavioral complexity, we measure the entropy of the distribution of actions performed over time and units of the same side (Fig. 3.b).
Visual Diversity and Quality-Weighted Visual Diversity
Coverage and QD-Score are classic measures of QD algorithms but are not easily computed using the archive in a large behavior space. As a proxy, we compute a single 2D projection using PCA (the two components capturing 2222% and 1212% of the explained variance) for all the behaviors of the intergenerational tournaments for all replications (Fig. 4), discretize this space with a $100\times 100$ grid, and compute the corresponding elites. Coverage (Fig. 3.c) corresponds to the proportion of non-empty cells, and QD-Score (Fig. 3.d) to the average fitness of the elites.
Quality
One limitation of computing the quality directly from the fitness in the QD-Score is that the fitness can be high because the opposite side is weak, not because the winning side is strong. To measure a less ambiguous quality of the elites found by each method, we follow Dharna et al., (2024) and select the best $10$ solutions from each side for each run of each variant, perform a round-robin tournament, and compute the ELO score (Elo,, 1978) (Fig. 5).
Open-endedness
Coverage measures the volume of novelty discovered by GAME with the VEM. One limitation is that this behavior is dependent on two solutions. Another limitation is that visual behavior may not capture the full range of novelty. We use another measure similar to PATA-EC proposed in Enhanced-POET. The idea is that a BT is different from another if, in a tournament, the ranking it generates in the opposition is different. Using the intergenerational tournament, we compute the ranking of all the BTs and count the number of new rankings per generation compared to the previous generations (Fig. 3.e). The idea is that an open-ended system will continue to create new challenges that require new solutions.
Results
Figure 4: PCA projections of the tournaments’ behaviors. GAME variants with a VEM show the most uniform coverage with less variance between replications.
Figure 5: Round-robin tournament ELO score between the $10$ best solutions of each replication. GAME-SO is significantly better than all variants.
Solution Size (Fig. 3.a)
Quality-only has a significantly higher increase in BT size. This can be explained by the lack of diversity, which focuses the optimization on a smaller set of solutions (one per task instead of $N_{cells}=25$), thus allowing more mutations to accumulate. Diversity-only has the second-largest increase in BT size, which can be explained by the absence of fitness regulation, leading to preserving “bad” nodes that would have been removed otherwise. Among all the GAME variants, GAME-SO shows a slightly higher increase in BT size.
Behavioral Complexity (Fig. 3.b)
Quality-only also has the highest increase in behavioral complexity, with the same explanation as before. What is more interesting is the presence of GAME-SO as a close second. Compared to GAME-MO, this suggests that minimizing BT size prunes necessary stepping stones that lead to more complex behaviors. Diversity-only is ranked fourth, showing that increased BT size does not necessarily lead to greater behavioral complexity.
Visual Diversity (Fig. 3.c and Fig. 4)
Quality-only and GAME-SO (no VEM) exhibit significantly less diversity than other variants, which show similar diversity levels, with Diversity-only being slightly superior, validating the use of a VEM as a behavior space.
Quality-Weighted Visual Diversity (Fig. 3.d)
Quality-only and GAME-SO (no VEM) show significantly worse QD-Scores (due to their poorer coverage). Diversity-only shows only a slightly worse QD-Score, suggesting that diversity alone is already a powerful optimizer, though the best QD-Score is achieved when also considering quality.
Quality (Fig. 5)
Comparing the best solutions’ quality, we observe that GAME-SO and some instances of Quality-only produce significantly better BTs. This suggests that increased behavioral complexity leads to significant performance improvements and that directly attempting to reduce solution size with a secondary objective may prevent the discovery of the highest-performing solutions. Quality-only has the largest variance between replications; this result suggests that diversity, quality, the VEM, and the bootstrap phase are all necessary to discover the best BTs.
Open-endedness (Fig. 3.e)
GAME-SO and GAME-MO show similar open-endedness, which slightly decreases with each iteration, finding 8080% new tasks in the 2nd and 3rd generations and declining to 6060% by the final generations. GAME-SO (no bootstrap) consistently generates 9090% novel BTs, suggesting this variant is the only one showing signs of true open-endedness throughout all $20$ generations, but also that bootstrapping may be detrimental to open-endedness. The other variants demonstrate less open-endedness, with Quality-only getting close to 0% novelty. This further demonstrates that diversity prevents getting trapped in local minima.
Two illumination takeaways
Figure 6: Global trends: Percentage of elites using (a) the Attack atomic with different targets and (b) the Go_to atomic for different thresholds through the generations. The solid line represents the median, and the shaded area shows the min and max of GAME-MO’s three replications.
Figure 7: Arms race example: Percentage of Red elites using the Attack atomic on the weakest enemy archer and Blue elites using the Heal atomic on random ally archers through the generations for GAME-SO’s three replications. For replication 22, as more Red elites targeted the Blue archers, Blue elites evolved to heal their archers.
GAME’s goal is to illuminate the solution space. We looked at the elites of each generation across GAME-MO’s three replications and examined the atomics they used to determine global trends. We discovered two interesting findings.
Regarding the Attack atomic (with any unit type), all replications converged to favor attacking either the farthest unit or a unit at random (Fig. 6.a). Attacking at random spreads damage and avoids wasting resources on overkills. Attacking the farthest unit is also beneficial because grenadiers deal area damage that can be avoided if they target a unit sufficiently far away.
Examining the Go_to atomics, all replications favor using a threshold distance equal to 5050% of the sight range while avoiding exact positioning, i.e., 0% (Fig. 6.b). This likely represents a balance between spending time moving precisely to the target position (i.e., 0%), which prevents units from performing other actions, and merely moving within sight of the target (i.e., 100100%), which limits visibility of what can be observed from the target position itself.
An example of arms race
We found an arms race in one of GAME-SO’s replications (Fig. 7). The Red elites increased their focus on the archers, leading to the Blue elites’ counter-strategy to focus their healing on their archers.
Case study: 2D Soft robots, Wrestling
Environment
Replicating the biological evolution feat of creating the diversity of living creatures’ morphology has been a long-standing research topic in the artificial life community (Langton,, 1997). One early work involves evolving articulated rigid-body creatures in an adversarial problem where the goal is for each creature to get as close as possible to a central cube (Sims,, 1994). To move closer to real creatures, some more recent works focus on soft robots with the problem of locomotion (Cheney et al.,, 2014; Kriegman et al.,, 2017; Cheney et al.,, 2018; Bhatia et al.,, 2021; Mertan and Cheney,, 2023). EvoGym (Bhatia et al.,, 2021) was proposed to facilitate the creation of environments for coevolving morphology and control of 2D soft robots.
While the coevolution of morphology and its control is an interesting and challenging topic, it is beyond the scope of this introductory GAME paper to explore the intersection of adversarial coevolution and body–brain coevolution. We instead focus on the adversarial coevolution of morphology in an adversarial environment, where the morphology passively provides actuation. We build a custom environment using EvoGym (Bhatia et al.,, 2021), which we refer to as Wrestling for 2D soft-robot simulation. We follow Cheney et al., (2014); Kriegman et al., (2017) by defining passive and active voxels that change surface following a sine wave pattern. We defined seven types of voxel: Empty (invisible); rigid (black); soft (gray); horizontal actuated in-phase (orange); vertical actuated in-phase (royalblue); horizontal actuated anti-phase (gold); vertical actuated anti-phase (skyblue).
We set the sine wave period to $12$ timesteps because it empirically results in most of the randomly sampled robots exhibiting non-stationary behaviors. The solution space is $5\times 5$ 2D soft robots, with the constraints that there is at least one actuated voxel and that all voxels are connected. See Fig. 8 for examples of different morphologies. The variation operator randomly chooses between adding, deleting, or mutating a voxel ($\frac{1}{3}$ for each) while respecting the constraints and does this $k$ times ($k=3$ in this paper).
In Wrestling, two robots start symmetrically on the edge of an arena of width 30 voxels. The fitness function is the percentage of timesteps the robot was closest to the center after $200$ timesteps. This leads each robot to reach the center as quickly as possible, push the other, and resist the other’s pushes. One specificity of this fitness function compared to the one used in Parabellum is that it is a relative measure of quality, i.e., given two competing blue and red solutions, $s_{blue}$ and $s_{red}$, $f(s_{blue})=1-f(s_{red})$. This means that even trivially bad solutions (e.g., standing morphologies) would lead to at least one of the two opposing fitnesses being positive. This prevents a direct evaluation of quality from the QD-Score and instead requires the use of a tournament or another measure of quality (e.g., the velocity of the solutions). This is interesting because a one-sided illumination would require already knowing a high-quality solution to avoid trivially overfitting a bad solution.
The behavior space is, similarly to Parabellum, the concatenation of the visual embedding of CLIP Radford et al., (2021) for five frames of the video. For speed optimization, we do not use EvoGym’s visualization but a reimplementation in JAX that directly creates RGB arrays for CLIP.
Results
Figure 8: Wrestling illumination with sampled snapshots. The central point cloud is the 2D PCA of the visual embedding of one illumination from GAME of the Wrestling adversarial problem (capturing 13.4% and 8.4% of the variance). The outer edge snapshots show examples of frames captured at the end of a diverse set of evaluations. The PCA captures slow-moving robots on the right and faster-moving robots on the left, each with a large variety of morphologies. See Figure_8.mp4 in the repository for a video of this figure.
Figure 9: Wrestling 2D PCA of the visual embedding with fitness. We concatenated all visual embedding behaviors and performed one PCA capturing 13.4% and 8.4% of the variance, and show the corresponding projection for GAME and Random (three replications). GAME leads to wider coverage than using the EvoGym random sampler. Wrestling’s fitness function measures the relative quality of solutions, so examining it does not directly allow comparison of the quality of the discovered solutions.
Figure 10: Wrestling visual diversity through generations. GAME leads to higher coverage of the visual behavior space. The solid line represents the median, and the shaded area shows the min and max of GAME and Random (three replications).
Figure 11: Wrestling quality of the best morphologies. ELO score computed with a tournament between the $10$ best morphologies of each side of each variant’s replications. GAME finds better morphologies than Random.
Comparison with a Random ablation.
To demonstrate the generality of GAME, we apply the same implementation to Wrestling as the one used on Parabellum with just a smaller budget of evaluations ($N_{budget}=$20\,000$\$, $N_{gen}=10$, and $N_{cells}=20$). We compare GAME against an ablation, Random, that uses EvoGym’s robot sampling function instead of applying the variation operator to an elite of the current generation. We performed $3$ replications for each and the same number of overall evaluations.
Fig. 9 shows the 2D PCA (capturing 13.4% and 8.4% of the variance) of the visual embedding. GAME leads to a larger coverage with a median of 44.2% [44.6, 39.7] against 28.9% [29.9, 27.3] for Random (Fig. 10) and higher quality when comparing the top-10 elites of each replication in a tournament; the lowest GAME ranking is higher than the highest Random ranking (Fig. 11). Fig. 9 does not allow us to compare the quality of the solutions due to Wrestling’s relative fitness function. To investigate the reason for the higher quality in the tournament, Fig. 12 shows the same PCA projection with the average velocity of the two soft robots. We can see that GAME leads to finding morphologies with higher velocities than the Random ablation. The main component of the PCA of the visual embedding captures slower robots on the right side, which can be visualized in Fig. 8 (and the supplementary video), with faster robots on the left side and slower-moving robots on the right side that do not reach the center of the arena.
Figure 12: Wrestling 2D PCA of the visual embedding with velocity. Same PCA projection but showing the average velocity of the two soft robots. GAME leads to discovering robot morphologies with higher velocities.
Lineage visualization.
GAME keeps track of each ancestor of a solution. Fig. 13 shows the full lineage of the best Red and Blue solutions of the inter-variant tournament. Because we mutate at most $k=3$, the solutions change smoothly. In comparison, using a higher $k$ or a crossover could lead to abrupt changes. Still, the initial and final morphologies have very few voxels in common. Future work could focus on using an indirect encoding (Cheney et al.,, 2014) to allow a crossover and boost the variation operator’s ability to explore the search space.
Figure 13: Ancestry of the best morphologies in the Wrestling tournament. GAME keeps track of each solution’s genealogy through the generations and evaluations. (a) Red. (b) Blue. The variation operator mutates at most $k=3$ voxels, leading to smooth transitions from one morphology to another.
Morphological speciation.
Figure 14: Morphological species through generations. (a–b) Clustering of the different morphologies into species from all generations using k-modes and UMAP for visualization. The numbers point to the selected elite of each cluster for each generation. (c–d) Average ELO score of each species through generations. GAME discovers all species in the first generation and then improves their average fitness.
To study the evolution of different morphologies through generations, we perform a posteriori morphological speciation for one of GAME’s replications. We first aggregate all the elites’ morphologies from each generation and perform clustering using k-modes ($k=5$) to identify different morphological species. Fig. 14.a–b show the UMAP projection of the clustering. Then, for each cluster at each generation, we compute the average ELO score in an intergenerational tournament. Fig. 14.c–d show the variation of performance through the generations for both sides, with a snapshot of the best morphology for each cluster and generation.
First, all the species are discovered in the first generation, even though the clustering is performed a posteriori on all generations. This means that GAME does not discover a significant new morphology after the first generation. This can be explained by two factors: (1) GAME is a QD algorithm, so it searches for diversity from the start, and (2) this specific problem is not open-ended.
Second, for most morphological species, the average ELO score improves over generations, demonstrating GAME’s ability to pressure for improved solution quality. We can notice that one Blue species disappears in the last generation. This discrepancy can be attributed to the difference between morphological diversity and the visual embedding diversity optimized by GAME, which captures other elements, such as velocity (due to the use of five frames), and the interaction between the two soft robots.
Case study: Deck building, Hearthbreaker
Figure 15: Hearthbreaker intergenerational tournament. ELO score of the elites from GAME execution on the five pairs of classes in Hearthbreaker in an intergenerational tournament for each pair. GAME leads to an overall improvement of the elites’ quality at each generation for both sides and all five pairs of classes.
Figure 16: Hearthbreaker archive comparison between GAME and ME. GAME leads to better top-50 elites than ME for 7 out of 10 comparisons and is not significantly better for the other 3. (Significance levels of the Mann–Whitney U test: $^{\ast}=p<0.05$, $^{\ast\ast}=p<0.01$, $^{\ast\ast\ast}=p<0.001$, and ns = not significant.)
Figure 17: Hearthbreaker measures comparison between GAME and ME. ME leads to denser and larger coverage due to focusing on only one task, while GAME illuminates 50 tasks simultaneously. Still, GAME finds solutions with similar quality to ME, even though it was not explicitly searching for solutions against the starting decks.
Warrior VS Warlock Rogue VS Paladin Hunter VS Druid Shaman VS Mage Mage VS Priest
(a) (b) (c) (d) (e) (f) (g) (h) (i) (j)
Coverage GAME 33.8% 34.3% 32.3% 33.5% 35.5% 30.0% 36.3% 27.3% 38.3% 33.0%
ME 66.0% 70.8% 65.3% 61.5% 66.3% 69.3% 64.0% 63.8% 66.8% 71.5%
QD-Score GAME 6.3 4.7 7.1 7.1 5.6 3.4 8.2 5.8 5.6 7.5
ME 12.5 9.1 13.3 13.1 10.8 6.2 14.9 12.1 9.3 16.4
Best Elite GAME 24.6 23.7 27.6 27.3 27.1 21.8 26.6 27.5 24.1 27.2
ME 26.1 23.4 27.9 27.1 26.7 22.6 28 28.0 25.3 28.2
Table 1: Illumination comparison between GAME and ME in Hearthbreaker against the starting deck. Coverage, QD-Score, and best-elite fitness of each execution of GAME and ME against the corresponding starting deck (Fig. 17). ME leads to higher coverage and QD-Score but similar highest fitness.
Environment
As a third adversarial problem, we aimed to compare GAME with a classic MAP-Elites approach that fixes one side of the adversarial problem, as in Fontaine et al., (2019). We use a Python simulator of the Hearthstone digital collectible card game (Blizzard Entertainment,, 2014), Hearthbreaker Daniel et al., (2014), for easy interfacing with our implementation of GAME.
Hearthstone is a two-phase card game. In the first phase, deck building, the player chooses a class (among the $9$ available) and creates a deck of $30$ cards. Then, in the second phase, deck playing, the player picks their deck of cards and plays a game against another player with their own class and deck of cards. The goal is to set the opponent’s life from $30$ to 0. In this case study, similarly to Fontaine et al., (2019), we only consider the deck building phase and use the Hearthbreaker Trade Agent heuristic to play the deck. This heuristic is a greedy, handcrafted policy that maximizes the loss of minions for the opponent and subsequently maximizes the damage to the opponent’s health.
We attempted to follow the method described in the original paper as closely as possible and did not use the VEM contribution of this paper as the behavior space. The solution space consists of a set of $30$ cards, comprising both class-specific cards and common cards (with between 250 and 300 available cards for each class in Hearthbreaker), with no more than two copies of the same card. Available cards may differ between the different versions of Hearthstone. We used the same variation operator, i.e., we pick one elite at random, pick $k$ from a geometric distribution with $P(k=1)=0.5$, and then randomly replace $k$ cards from the elite deck using a uniform distribution over the valid cards. The behavior space is two-dimensional: the deck’s mana cost mean and standard deviation. We use the same unstructured archive, as it can handle both small and large spaces without the same value of hyperparameter, i.e., the number of cells in the archive. The fitness function is the difference in life between the player and the opponent. As the game ends when one player reaches 0, the fitness function is positive when the player wins (and is bounded by the starting health of $30$) and negative when the player loses (and is bounded by 30-30). As the game is stochastic by nature due to the shuffling of the cards, the fitness is averaged over $n=50$ duels with fixed seeds, with half of them starting with the player and the other half with the opponent.
Results
We compare GAME against an ablation, ME, corresponding to MAP-Elites, i.e., only one generation and one task, with the same total budget of 40 00040\,000 evaluations per side and number of cells in the archive. GAME is set with $N_{gen}=8$, $N_{budget}=$10\,000$\$, $N_{tasks}=50$, and $N_{cells}=$20$\$ and ME with $N_{gen}=1$, $N_{budget}=$40\,000$\$, $N_{tasks}=1$, and $N_{cells}=$1\,000$\$. We chose five pairs of classes that are empirically known to be balanced: Warrior and Warlock, Rogue and Paladin, Hunter and Druid, Shaman and Mage, and Mage and Priest. ME independently evolves the deck of one side against the fixed starting deck of the other side (similarly to Fontaine et al., (2019)) and is thus executed once for each side. In contrast, GAME coevolves both decks in one run. Fig. 15 shows the intergenerational ELO score of the five GAME executions on each pair of classes. We can see that for all instances, the elite’s average quality improves at each generation.
As GAME and ME do not use the same resolution for the archive, for a fairer comparison, we recomputed both variants’ archives in a 2D grid with a bin size of $0.1$, and performed a tournament between the new archives’ elites for each pair of classes. Fig. 16 shows the resulting ELO scores. We can see that GAME finds significantly better solutions for 7 out of 10 comparisons and non-significantly better solutions for the remaining three. A simple reason is that ME decks are only evolved to compete against the starting deck, whereas GAME coevolves a diversity of decks that prevent overfitting to a single one.
To compare the illumination, we evaluated each elite from GAME against the corresponding starting deck (the same used as a fixed opponent for ME) and compared the corresponding diversity and quality. Fig. 17 shows each archive, and Tab. 1 shows the corresponding Coverage, QD-Score, and best-elite fitness. GAME’s coverage is significantly worse than ME’s. The reason is that it splits the 1 0001\,000 cells into $50$ tasks, thus leaving only $20$ cells for diversity, while ME can allocate the 1 0001\,000 cells for diversity. This becomes apparent in this case study because we use a much smaller 2D behavior space, rather than a visual embedding. It is thus much easier to cover the entire reachable behavior space. One limitation of GAME is the independence of diversity in each task, which leads to storing elites with similar behaviors across multiple tasks, thereby reducing the overall diversity of the archive. Still, even though GAME’s elites are not evolved against the starting decks like those from ME, they show similar performance against them.
Discussion and Future Work
Multi-agent battle game: Parabellum
We hypothesized that minimizing the BT size as a secondary objective would not impact quality and would lead to smaller, more interpretable BTs. However, the results indicate it prunes the neutral mutations that seem to be essential stepping stones for high-performing solutions. This supports the neutral theory of molecular evolution (Kimura,, 1979), which states that most variation at the molecular level is neutral yet leads to the evolution of complex organisms in nature.
We also hypothesized that bootstrapping each generation using solutions from previous ones would accelerate the search compared to starting from scratch. The results confirm this hypothesis. Nonetheless, GAME-SO (no bootstrap) is the only variant demonstrating constant generation of new solutions throughout all $20$ generations. This phenomenon could be related to extinction events (Lehman and Miikkulainen,, 2015) and warrants further investigation, as it currently does not yield the best diversity or quality.
One limitation of our current evaluation of GAME is that Parabellum is a symmetrical game that may not present imbalance issues between sides. For example, POET (Wang et al.,, 2019), which coevolves controllers and environments, must rely on a selection of environments that are neither too easy nor too difficult for the current population of controllers. Future work should examine the application of GAME in asymmetric adversarial problems.
One limitation of this case study is the use of BTs as controllers. They are inherently interpretable but require handcrafted atomics that limit the possible behaviors. An interesting future direction is to utilize end-to-end neural networks that map observations to actions, potentially employing neuroevolution Stanley and Miikkulainen, (2002), to move toward a more open-ended search space.
2D Soft robots: Wrestling
The main limitation of this case study is the use of passive actuation to control the robots, which, in particular, prevents the robots from exhibiting reactive behavior based on sensor inputs, a key component of intelligent behaviors (Brooks,, 1991; Pfeifer and Bongard,, 2006). Future work should focus on incorporating body–brain coevolution with GAME to move toward a truly open-ended search space for artificial creatures.
Deck building: Hearthbreaker
There are two limitations of the current approach. First, the behavior space may be too simple and is independent of the duels (as it is purely computed from the deck and not influenced by how the deck is played), which limits the possible diversity of decks. Second, we use a greedy heuristic that ignores the combo aspect of the game and focuses only on building decks with many minions (all decks evolved with GAME have at least 86% (26 out of 30) minion cards, with a median of 97% (29 out of 30). In contrast, minions correspond to 65% of the available cards, which would correspond to 20 out of 30 cards in the deck. Future work should focus on using a more advanced agent, such as MCTS (Santos et al.,, 2017), to explore the game’s combo aspect and thus significantly increase the open-endedness of this adversarial problem.
Conclusion
We present GAME, a coevolutionary QD algorithm for illuminating adversarial problems. Combined with a vision embedding model as a domain-agnostic behavior space, GAME requires only videos instead of handcrafted behavior descriptors. We demonstrate GAME’s ability to illuminate three different adversarial problems and validate all its components through ablation studies in a multi-agent battle game. We show its generality by using the same code (GAME + VEM) when evolving both behavior trees for a multi-agent battle game and soft-robot morphologies in a wrestling environment. The deck building illumination also demonstrates its strength in fostering quality, while revealing limitations in covering small behavior spaces due to its multi-task formulation. This work is limited by the non-open-ended nature of the different search spaces, which prevents GAME from generating open-ended illumination. Future work should focus on applying GAME to open-ended search spaces, paving the way toward a better understanding of the emergence of open-ended adversarial coevolution.
Acknowledgments
Funded by the armasuisse S+T project F00-007.